# spacing=nonfrench

### v1

@@ 4 Threat Model
Our threat model describes online users using GenAI browser assistants to personalize their browsing experience. The browser extension of the assistant installed onto the user’s browser is assumed to be the adversary while the user is considered the victim. The adversary operates in context of an extension providing it access to everything that the user does on the web. The primary goal of the adversary is to provide personalized responses to user’s queries. However, in order to achieve its primary goal, the adversary is assumed to engage in tracking victim’s browsing activities – by leveraging its presence as an extension. Thus user tracking is considered the secondary (or implicit) goal of the adversary. Adversary may share the collected information from the victim’s browser – either to its own servers or with third-party servers.
Figure 1 depicts our threat model. The victim can interact with the GenAI browser assistants in a variety of different ways – for example, making a Google search, summarizing the browsed pages, chatting with the assistant about the page, etc. Any such activity that involve a user’s interaction with the GenAI assistant is assumed to be a user’s query. Upon receiving a user query, assistant sends it along with metadata to its own server. The metadata may constitute user-, device-, webpage-, or browsing-specific information that serves as a “context” to generate a personalized response. GenAI assistant’s server may either generate a personalized response in-house or share the user’s query and context to an open-source LLM model using their API to generate a response. Finally, the personalized response is sent to the user’s browser and displayed to the victim
The GenAI browser assistant may also set and share first-party cookies from the victim’s browser to its own server. These cookies could further be shared with respective tracking platforms to either learn more about the user (i.e., analytics) or create custom audiences to re-target them with personalized ads on those third-party platforms. To fulfill its primary goal, adversary is assumed to continuously build and refine a user’s profile using the collected context. We consider LLM model’s server or third-party trackers to be beneficiaries as they may benefit from the data provided by the GenAI-browser assistant but are not actively involved in explicitly collecting information related to the victim.
5 Methodology
In this section we introduce our novel auditing framework for GenAI browser assistants to audit user tracking, profiling, and personalization as depicted in Figure 2. We make our crawling and analysis framework available at https://anonymous.4open.science/r/gen-ai-privacy-audit-1C72.
5.1 Selection of GenAI Browser Assistants
@@
Sider: ChatGPT Sidebar 4M [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] sider Automatic Server-side PII WC 4.35.0
Monica - Your AI Copilot 2M [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] gpt-4o-mini Mixed Server-side PII UA PC FI 7.6.0
@@
MaxAI: Chat with Webpage 800K [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] gpt-4o-mini Manual Server-side PII UA 6.7.1
@@
Wiseone - AI Copilot 90K [Uncaptioned image][Uncaptioned image] gpt-4o Manual Server-side PII WC 1.7.2
@@
Among the three categories of GenAI browser assistants described in Section 2, we focus on search-based assistants. This is due to several reasons: (1) Querying a search engine is relevant to querying an LLM model, allowing a natural integration of LLM-generated responses to a user’s online searches. (2) It has been argued that search engines and LLMs would co-exist due to their unique individual capabilities to incorporate recent information and generate context-aware responses, respectively [32]. (3) Although popularly referred to as ‘search’ based assistants, their capabilities are not restricted to search-only and they provide functionalities that allow their usage on any webpage as opposed to other categories of assistants that have limited scope and functionalities. (4) Combination of continued access to a user’s entire web browsing alongside generative capabilities raises privacy concerns, making them perfect candidates for our audit.
Our audit focuses on the most widely used web browser – Google Chrome. We survey various extensions from the Chrome Web Store to identify "AI-based search assistants" that offer the following three fundamental capabilities. (1) direct integration with browser search (i.e., Google search) via a sidebar response; (2) interactive sidebar functionality to chat with any webpage; and (3) ability to summarize webpage content. We select the top 10 most popular extensions based on number of installations that satisfy these requirements. The details about the selected assistants are shown in Table 1.
@@ 5.2 Crawling Infrastructure
To audit data collection and tracking practices of these browser assistants, it is important to examine network traffic activity emerging from the browser with the installed extension. We use a dedicated laptop to carry out our auditing experiments. The laptop is connected to a wireless Wi-Fi network to access the Internet and has Chrome browser installed. To intercept and log the decrypted network traffic while browsing the web, we install and configure Mitmproxy [39] – an open-source HTTPS proxy – on the device following official guidelines [19]. During the proxy configuration phase, Mitmproxy certificate is imported into Chrome’s Trusted Root Certification Authorities, allowing Mitmproxy to intercept and decrypt HTTPS traffic without triggering security warnings. The proxy server in this set-up acts as a man-in-the-middle between the laptop and the rest of the Internet as depicted in Figure 2, allowing us to capture all incoming as well as outgoing real-time web traffic while conducting our experiments – including requests, responses, payloads, and stored identifiers like cookies.
@@
To conduct different experiments described in Sections 5.3 and 5.4 – each time, we first initialize a new Mitmproxy instance on localhost with mitmweb using command line interface (CLI). Next, we launch Chrome browser from CLI by specifying a new custom directory for user data to open a fresh browser instance with a new profile each time (Step ). This ensures a clean browsing session with no past history associated with the profile, resulting in independence across different experiments. Next, we visit the Chrome web store, search for the browser assistant, and install its extension into the browser instance (Step ). All assistants in our study, except Perplexity, require a user login to use it. Perplexity can be used with or without logging-into the account. To keep the profiles isolated across different experimental scenarios, we sign-up for a new user account each time using a distinct temporary email provided by various online services (Step ). However, some assistants doesn’t allow the use of a temporary email address – restricting it to either a Google or an Apple account. In such cases, we re-use an experimental Google account as Google prohibits creation of unlimited accounts tied to the same phone number. To avoid contamination across different experimental scenarios audited for a given extension, before the start of each experiment, we manually open a fresh browser instance to delete all Google activity and account-associated history. We also delete chat history and account-associated memory with the browser assistant’s account. We acknowledge that the browser assistant may store a server-side mapping of the user’s account-related activity data – which may or may not be deleted when the user performs client-side deletion. This is a limitation of our methodology. Next, we conduct the experiments described in the following sections. Finally, at the end of each experiment, web traffic from the entire browsing session is stored into a .flow file. We describe analysis of .flow files in Section 5.5.
@@ 5.3 Auditing User Tracking
When navigating through the web with an installed GenAI browser assistant, a user may visit a mix of public as well as private spaces online. In these spaces, a user may use the assistant without realizing the consequences of doing so. For instance, a user browsing through their health records may inadvertently end-up sharing sensitive information with the use of browser assistant. It is important to understand if the assistants are responsible and private with in-built safeguards to avoid exfiltration of user’s details or do they freely collect, share, and store information about the user at all times.
Experimentation. To understand implicit data collection and sharing, we identify 20 commonly-browsed website categories – 10 public spaces and 10 private spaces as listed in Table 6. For each category, we select one popular website to experiment with. All websites in private space are authenticated and contain personal or sensitive information that a user would not want to be collected, stored or shared. We experiment each website for every extension, totaling to a set of 200 experiments. Using our crawling infrastructure, we visit the website after logging-into the browser assistant’s account. Next, if it is a private space website, we login to the test website using (our own) personal account. If applicable, we navigate through the website by clicking on specific links that land us to the webpage containing private information about the user. For example, we navigate to ‘Medical Records’ tab and click on a specific visit record to open a detailed medical record associated with some visit. To avoid inconsistencies and ensure comparability, we always perform the same set of actions (if any) for a given website across different experiments. Once the website is loaded and navigations are complete, we invoke the sidebar interface and perform summarization of the webpage using its ‘summary’ feature (Step ). Finally, we ask a follow-up question (see Table 6 in appendix) to explicitly seek some information displayed on the webpage (Step ). This is to understand if the assistant recognizes the page content to be sensitive, problematic, copyrighted, or personal to avoid providing the requested information or not. Responses from the browser assistants as well as .flow files are stored for analysis described in Section 5.5.
@@ 5.4 Auditing Profiling and Personalization
As defined in Section 1, profiling involves collecting data pertaining to the users to draw inferences about them. However, using such a holistic view inferred from the profiled information to create individualized experience is personalization. It is important to understand if GenAI browser assistants associate the collected information about the user to their profile that it can recall to answer specific questions by the user, or to personalize its responses.
@@
Prompting Framework. To evaluate profiling and personalization in GenAI browser assistants with respect to RQ3, we propose a novel prompting framework as depicted in Figure 2. The profile we develop represents – “A rich millennial male from southern California who is interested in equestrian activities”. We formulate 5 Train-Test prompt pairs – each explicitly leaking some attribute about the user in the train prompt and then testing for profiling based on the leaked information via a test prompt. The user’s attribute that we leak and test for in chronological order of prompts include location, age, gender, wealth, and interest. After the last test prompt, we re-ask all the 5 test prompts combined together as a single “profiling prompt” to the GenAI browser assistant. This is followed by “personalization prompt” where we simply ask the assistant to suggest Top 3 activities that the user would likely include in their itinerary of a vacation, based on what it has learnt about the user. The profiling and personalization prompts are first asked in the same tab where rest of the training or testing of individual attributes occur to understand its in-context behaviour. Next, we also open a new browser tab within the same session and re-ask them to understand its out-of-context behaviour. Except personalization and train prompts, we condition all responses to 5 test prompts and the profiling prompt to only output a binary response – "Yes" or "No" to avoid any subjectivity. This is achieved by adding a meta (system) prompt. Personalization prompt is reported to personalize responses (i.e., "Yes") if one of the suggested activities clearly lists atleast one ‘equestrian activity’ (i.e., the user’s leaked interest). We explicitly choose a niche interest for our profile to avoid matches with generic activities that an LLM model might suggest such as ‘hiking’ for instance. Section 8.2 in appendix lists all prompts used in this study.
@@
Experimentation. Amongst different features provided by search-based GenAI browser assistants, we specifically test three most useful features that were observed to be common across all 10 extensions – search-based integration, webpage chat, and webpage summary. These are studied under the following four scenarios for each extension (totaling to 40 experiments) as described in Figure 2:
@@
Search. In search scenario, train prompts are entered into Google search as a normal search query. When Google search is performed, it may result in an automatic response generation from the GenAI assistant as well, which is displayed to the user as a sidebar response. In cases where automatic responses were not generated, the train prompt is explicitly asked to the GenAI assistant in the sidebar chat. Next, the test prompt corresponding to the train prompt is asked as a follow-up question in the sidebar chat. The same steps are repeated for each train-test prompt pair within the same browser tab. At the end, profiling and personalization prompts are asked both – in-context and out-of-context.
Browse. In this scenario, no interaction occurs during the training phase. Training phase involves browsing through 10 webpages – 2 pages per leaked attribute. These webpages (listed in Table 5 in appendix) are selected such that they leak different attributes about the user. Other than scrolling a page to explore its content, we also click on internal links to emulate normal browsing. The set of clicks are fixed and remains the same across different experiments involving the same website. Our hypothesis is that if GenAI browser assistant collects information about the user’s visit to different webpages, then it may use it to infer user’s attributes. The experiment ends by asking 5 test prompts along with profiling and personalization prompts sequentially within the same tab in the sidebar chat. Profiling and personalization prompt are also repeated out-of-context.
Summarize. We use “summarize webpage” feature of GenAI browser assistants to understand if usage of this feature aids the browser assistant in building a profile about the user and later on use it to personalize its responses. We visit the same 10 webpages as before, performing the same interactions as in ‘Browse’. Additionally, we also summarize each page. In this scenario, summarization acts as training. Finally, the same prompting as in ‘browse’ is performed.
We repeat each prompting experiment in our study 3 times and report the majority response to account for variability in the output due to probabilistic nature of GenAI systems.
@@ 5.5 Semi-automated Analysis
In this section, we discuss our approach to perform network traffic analysis based on the web traffic data collected in the form of .flow files during the experiments. We analyse .flow files corresponding to the most representative (i.e., majority) response. We use Python module mitmproxy.io to programmatically parse the network flows and extract details about the contacted endpoints, request and response headers, payloads, and responses. We identify each request to be first-party if its domain matches the domain of the browser assistant’s website, and third-party otherwise. A flow is also classified as either a foreground flow or a background flow. Background flows represent extension traffic emerging from the background service worker of the browser assistant. They are identified based on the value of a request header – origin – value of the form chrome-extension://<extensiod-id> represents traffic from an extension, where extension-id can be obtained from extension’s page on Chrome web store. The automated code extracts the flows from different experiments related to a given extension along with the above details in the form of a CSV file. We perform detailed analysis of these files to extract user, device, webpage or browsing related information from the payload.

### v2

@@ 4 Threat Model
Our threat model considers online users who install a GenAI browser assistant to personalize their browsing experience. In this threat model, the browser extension of the assistant is treated as the adversary while the user is the victim. The adversary operates as a browser extension, which provides it access to a user’s complete browsing activity. The primary goal of the adversary is to provide personalized responses to user queries. To achieve this, it is assumed to engage in user tracking by monitoring browsing activities and extracting contextual information. This user tracking is considered the secondary (or implicit) goal of the adversary. Adversary may share the collected information from the victim’s browser either to its own servers or to third-party servers.
Figure 1: Threat model of GenAI browser assistants
Figure 1 illustrates our threat model. The victim can interact with the GenAI browser assistants in different ways such as initiating a Google search, summarizing a web page, chatting about the web page’s content, etc. Any such activity that involves a user’s interaction with the GenAI assistant is assumed to be a user’s query. Upon receiving a user query, the assistant sends it along with metadata to its own backend server. The metadata may include user-, device-, webpage-, or browsing-specific information that serves as a context for generating a personalized response. The assistant’s server may generate the response internally or share the query and context to an external LLM API. The personalized response is sent back to the victim’s browser.
The assistant may also set and share cookies from the victim’s browser to its own server. These cookies could further be shared with third-party advertising or analytics platforms to support analytics or ad targeting, such as building custom audiences for personalized advertising. To fulfill its primary goal, the adversary is assumed to continuously build and refine a profile of the user based on accumulated context over time. We consider the LLM provider and third-party trackers as beneficiaries of the data collected and shared by the assistant. Although they may gain this information, we do not assume them to be directly involved in data collection.
@@
Sider: ChatGPT Sidebar 5M [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] sider Automatic Server-side PII WC 4.35.0
Monica - Your AI Copilot 3M [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] gpt-4o-mini Mixed Server-side PII UA PC FI 7.6.0
@@
MaxAI: Chat with Webpage 900K [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] gpt-4o-mini Manual Server-side PII UA 6.7.1
@@
5 Methodology
In this section, we introduce our auditing framework for GenAI browser assistants, designed to evaluate user tracking, profiling, and personalization, as shown in Figure 2. We make our crawling and analysis framework available at https://doi.org/10.5281/zenodo.15618960.
5.1 Selection of GenAI Browser Assistants
Among the three categories of GenAI browser assistants described in Section 2, we focus on search-based assistants. This is due to several reasons: (1) Querying a search engine is relevant to querying an LLM model, allowing a natural integration of LLM-generated responses to a user’s online searches. (2) It has been argued that search engines and LLMs would co-exist due to their unique individual capabilities to incorporate recent information and generate context-aware responses, respectively [30]. (3) Although commonly referred to as “search-based,” these assistants’ capabilities are not limited to search alone; they provide functionalities that allow their usage on any webpage as opposed to other categories of assistants that have limited scope and functionalities. (4) Combination of continued access to a user’s complete browsing activity alongside GenAI capabilities raises privacy concerns, making them ideal candidates for our audit.
Our audit focuses on the most widely used web browser, Google Chrome. We survey various extensions from the Chrome Web Store to identify “AI-based search assistants" that offer the following three core capabilities. (1) direct integration with search engines (e.g., Google search) via a sidebar response; (2) interactive sidebar functionality to chat with any webpage; and (3) the ability to summarize webpage content. We select nine of the most popular extensions that meet these criteria, based on installation numbers. Details about the selected assistants are provided in Table 1.
@@ 5.2 Crawling Infrastructure
To audit data collection and tracking practices of GenAI browser assistants, it is important to examine network traffic activity generated by the browser with the extension installed. We use a dedicated laptop to conduct our experiments. The laptop is connected to the Internet via a Wi-Fi network and has Chrome browser installed. To intercept and log decrypted network traffic while browsing, we install and configure Mitmproxy [27] – an open-source HTTPS proxy – on the device following the official guidelines. During the proxy configuration phase, Mitmproxy certificate is imported into Chrome’s Trusted Root Certification Authorities, allowing Mitmproxy to intercept and decrypt HTTPS traffic without triggering security warnings. The proxy server acts as a man-in-the-middle between the laptop and the Internet, as shown in Figure 2. This allows us to capture all incoming and outgoing real-time web traffic while conducting our experiments, including requests, responses, payloads, and browser storage.
@@
To conduct the experiments described in Sections 5.3 and 5.4, we initialize a new Mitmproxy instance on localhost with mitmweb using command line interface (CLI). Next, we launch Chrome browser by specifying a new custom directory for user data to open a fresh browser instance with a new profile each time (Step ). This ensures a clean browsing session with no past history associated with the profile, providing independence across experiments. Next, we visit the Chrome web store, search for the browser assistant, and install its extension into the browser instance (Step ). All assistants in our study, except Perplexity, require a user login. Perplexity can be used with or without logging-into an account. To keep the profiles isolated across experiments, we sign-up for a new user account each time using a distinct temporary email provided by various online services (Step ). However, some assistants do not allow the use of a temporary email address and instead require authentication via Google or Apple accounts. In such cases, we reuse a designated Google account, since Google restricts the number of accounts tied to a single phone number. To avoid contamination across different experiments for a given extension, before the start of each experiment, we manually open a fresh browser instance and delete all Google activity and browsing history. We also delete the assistant’s chat history and any account-associated memory. We acknowledge that the browser assistant may store a server-side mapping of the user’s account-related activity – which may or may not be deleted when the user performs client-side deletion. This is a limitation of our methodology. After setup, we conduct the experiments described in the following sections. At the end of each experiment, web traffic from the entire browsing session is stored in a .flow file. We describe analysis of .flow files in Section 5.5.
@@ 5.3 Auditing User Tracking
When navigating through the web with an installed GenAI browser assistant, a user may visit a mix of public and private online spaces. In these spaces, a user may use the assistant without realizing the consequences of doing so. For instance, a user browsing through their health records may inadvertently share sensitive information due to the use of the browser assistant. It is important to understand if the assistants have built-in safeguards to prevent collection and exfiltration of user’s sensitive information or do they freely collect and share such information at all times.
To understand such implicit data collection and sharing, we identify 20 commonly-browsed website categories: 10 public spaces and 10 private spaces as listed in Table 6. For each category, we select one popular website to experiment with. Websites classified as private require user authentication and comprises personal or sensitive information that a user would not want to be collected or shared. We test each website with every extension, resulting in a total of 180 experiments (20 websites x 9 extensions). Using our crawling infrastructure, we visit the website after logging-into the browser assistant’s account. Next, if it is a private space website, we log in to the test website using (our own) personal account. If applicable, we navigate through the website by clicking on specific links that lead us to the webpage containing private information about the user. For example, we navigate to ‘Medical Records’ tab and click on a specific visit to open a detailed medical record associated with some visit. To avoid inconsistencies and ensure comparability, we always perform the same set of actions (if any) for a given website across different experiments. Once the website is loaded and navigations are complete, we invoke the sidebar interface and perform summarization of the webpage using its ‘summary’ feature (Step ). Finally, we ask a follow-up questions (see Table 6 in appendix) to explicitly request information displayed on the webpage (Step ). This is to understand whether the assistant recognizes the page content to be sensitive, problematic, copyrighted, or personal to avoid providing the requested information or not. Responses from the browser assistants as well as .flow files are stored for further analysis.
@@ 5.4 Auditing Profiling and Personalization
Profiling refers to the process of collecting user data to infer attributes such as demographics, preferences, or interests. Personalization builds on profiling by using these inferences to tailor responses or experiences to the individual user. It is important to understand whether GenAI browser assistants associate the collected information with the user’s profile to answer specific questions or personalize their responses
@@
Prompting Framework. To evaluate profiling and personalization in GenAI browser assistants with respect to RQ3, we propose a novel prompting framework as depicted in Figure 2. The profile we develop represents – “A rich millennial male from southern California who is interested in equestrian activities”. We formulate 5 Train-Test prompt pairs – each explicitly leaking some attribute about the user in the train prompt and then testing for profiling based on the leaked information via a test prompt. The user’s attribute that we leak and test for in chronological order of prompts include location, age, gender, wealth, and interest. After the last test prompt, we re-ask all the 5 test prompts combined together as a single profiling prompt for the GenAI browser assistant. This is followed by “personalization prompt” where we simply ask the assistant to suggest Top 3 activities that the user would likely include in their itinerary of a vacation, based on what it has learnt about the user. The profiling and personalization prompts are first asked in the same tab where rest of the training or testing of individual attributes occur to understand its in-context behaviour. Next, we also open a new browser tab within the same session and re-ask them to understand its out-of-context behaviour. Except personalization and train prompts, we condition all responses to 5 test prompts and the profiling prompt to only output a binary response – "Yes" or "No" to avoid any subjectivity. This is achieved by adding a meta (system) prompt. Personalization prompt is reported to personalize responses (i.e., "Yes") if one of the suggested activities clearly lists atleast one ‘equestrian activity’ (i.e., the user’s leaked interest).
Selection and Ordering of Prompts. We explicitly choose a niche interest for our profile to avoid matches with generic activities that an LLM model might otherwise suggest such as ‘hiking’ for instance. This minimizes the risk of spurious correlations between the leaked interest and the personalized responses. User descriptions of each prompt were crafted in accordance with the attribute we wanted to test for. To test for a new attribute, future work can easily generate a description that leaks the attribute by asking an LLM model and providing our prompts as a context. Additionally, each train prompt targets a distinct user attribute (e.g., location, age). Thus, we do not expect any meaningful difference from either the specific ordering or the prompt used. Section 8.1 in appendix lists all prompts used in this study.
@@
Experimentation. Amongst different features provided by search-based GenAI browser assistants, we specifically test three most useful features that were observed to be common across all 9 extensions – search-based integration, webpage chat, and webpage summary. These are studied under the following four scenarios for each extension (totaling to 36 experiments) as described in Figure 2:
@@
Search. In search scenario, train prompts are entered into Google search as a normal search query. When Google search is performed, it may result in an automatic response generation from the GenAI assistant, which is displayed to the user as a sidebar response. In cases where automatic responses are not generated, the train prompt is explicitly asked to the GenAI assistant in the sidebar chat. Next, the test prompt corresponding to a train prompt is asked as a follow-up question in the sidebar chat. The same steps are repeated for each train-test prompt pair within the same browser tab. At the end, profiling and personalization prompts are asked both – in-context and out-of-context.
Browse. In this scenario, no interaction occurs during the training phase. Training phase involves browsing through 10 webpages – 2 pages per leaked attribute. These webpages (listed in Table 5 in appendix) are selected from the top-ranking Google search results of the specific attribute descriptions leaked in the finalized prompts. We manually confirmed that the website content explicitly reveals the given attribute. Other than scrolling a page to explore its content, we also click on internal links to emulate normal browsing. The set of clicks are fixed and remains the same across different experiments involving the same website. Our hypothesis is that if GenAI browser assistant collects information about the user’s visit to different webpages, then it may use it to infer user’s attributes. The experiment ends by asking 5 test prompts along with profiling and personalization prompts sequentially within the same tab in the sidebar chat. Profiling and personalization prompt are also repeated out-of-context.
Summarize. We use the summarize webpage feature of GenAI browser assistants to understand if usage of this feature aids the browser assistant in building a profile about the user and later on use it to personalize its responses. We visit the same 10 webpages as before, performing the same interactions as in Browse. We also summarize each page. In this scenario, summarization acts as training. Finally, the same prompting as in Browse is performed.
We repeat each prompting experiment in our study 15 times and report the majority response to account for variability in the output due to probabilistic nature of GenAI systems.
@@ 5.5 Semi-automated Analysis
In this section, we discuss our approach to perform network traffic analysis based on the web traffic data collected in the form of .flow files during the experiments. We analyze the .flow files corresponding to the majority response across repeated runs. We use Python module mitmproxy.io to programmatically parse the network flows and extract details about the contacted endpoints, request and response headers, payloads, and responses. We identify each request to be first-party if its domain matches the domain of the browser assistant’s website, and third-party otherwise. To avoid any misclassifications due to organizational relatedness of different domains, we further use DuckDuckGo’s entity list that maps domains to their parent organization. If two domains were found to belong to the same organization, and one was initially classified as first-party while the other as third-party, we reclassified the latter as first-party. A flow is also classified as either a foreground flow or a background flow. Background flows represent extension traffic emerging from the background service worker of the browser assistant. They are identified based on the value of a request header – origin. A value of the form chrome-extension://<extension-id> represents traffic from an extension, where extension-id can be obtained from extension’s page on Chrome web store. An automated script extracts the flows from different experiments related to a given extension along with the above details in the form of a CSV file. We perform detailed analysis of these files to extract user, device, webpage, or browsing related information from the payload.